# Bundle Neural Networks for message diffusion on graphs

### v1

@@ 6.2 Real-world tasks
We evaluate BuNNs on the Long Range Graph Benchmark [32] and the heterophilic tasks from Platonov et al. [33] and. In our experiments, we report results both for the BuNN model described in Section 3 and a slight variation inspired by MixHop GNNs [34], which we call BuNN-Hop. We found this variation to perform exceptionally well on some real-world tasks. We provide the implementation details in the Appendix (Section E).
Table 3: Results for the heterophilic tasks. Accuracy is reported for roman-empire and amazon-ratings, and ROC AUC is reported for minesweeper, tolokers, and questions. Best results are denoted by First and Second. Asterisk denotes that some runs ran out of memory on an NVIDIA A10 GPU (24 GB).
@@
BuNN $91.75\pm 0.39$ $53.74\pm 0.51$ $98.99\pm 0.16$ $84.78\pm 0.80$ $78.75\pm 1.09$
BuNN-Hop $89.63\pm 0.59$ $52.01\pm 0.45$ $97.91\pm 0.30$ $84.33\pm 0.77$ $78.68\pm 1.04$
Heterophilic datasets. As we have shown in Section 4, BuNNs are provably capable of avoiding over-smoothing. It is, therefore, natural to test how BuNN performs on heterophilic graphs where over-smoothing is recognized as an important limitation (see, e.g., Yan et al. [35]). We follow their methodology to evaluate BuNN on the $5$ heterophilic tasks proposed in Platonov et al. [33]. We run the models with $10$ different seeds and report the mean and standard deviation of the test accuracy for roman-empire and amazon-ratings, and mean and standard deviation test ROC AUC for minesweeper, tolokers, and questions. We use the classical baselines from Platonov et al. [33] and provide the hyper-parameters in the Appendix (Section F.2).
We report the results in Table 3. BuNN achieves the best score on all tasks, with an average relative improvement of $4.4\%$. Its score on minesweeper is particularly impressive, which is significantly ahead of the rest and for which BuNN solves the task perfectly. We found that the optimal value of $t$ over our grid search varies across datasets, being $1$ for amazon-ratings and $100$ for roman-empire. BuNN-Hop also performs similarly but tends slightly inferior to BuNN on $3$ of the datasets and inferior on $2$. BuNN consistently outperforms the sheaf-based model NSD. Such results showcase the strong modeling capacity of BuNN in heterophilic settings.
Long Range Graph Benchmark. In Section 4.2, we show that BuNNs have desirable properties when it comes to over-squashing and modeling long-range interactions. To verify such claims empirically, we evaluate BuNN on tasks from the Long Range Graph Benchmark (LRGB) [32]. We consider the Peptides dataset consisting of 15 5351553515\,53515 535 graphs which come with two associated graph-level tasks, Peptides-func and Peptides-struct.
Table 4: Results for the Peptides-struct and Peptides-func tasks from the Long Range Graph Benchmark (results are ×$\times 100$ for clarity). The best result is bold.
Peptides-func Struct
Model Test AP $\uparrow$ Test MAE $\downarrow$
GCN $68.60\pm 0.50$ $24.60\pm 0.07$
GINE $66.21\pm 0.67$ $24.73\pm 0.17$
GatedGCN $67.65\pm 0.47$ $24.77\pm 0.09$
GPS $65.34\pm 0.91$ $25.09\pm 0.14$
DReW $71.50\pm 0.44$ $25.36\pm 0.15$
@@
BuNN $71.03\pm 0.22$ $24.98\pm 0.12$
BuNN-Hop $\mathbf{71.92}\pm 0.22$ $24.63\pm 0.25$
The graph classification task in Peptides-func is to predict the function of the peptide from $10$ classes, while the regression task in Peptides-struct is inferring the 3D properties of the peptides. In both cases, we follow the standard experimental setup detailed by Dwivedi et al. [32] alongside the updated suggestions from Tönshoff et al. [36]. The performance metric is Average Precision (AP) (higher is better) for Peptides-func and Mean Absolute Error (MAE) (lower is better) for Peptides-struct. We run each experiment on $4$ distinct seeds and report mean and standard deviation over these runs. Baseline models are taken from Tönshoff et al. [36] and include MPNNs, transformer models, and the current SOTA models [37, 38].
We report the results in Table 4. BuNNs achieve, to the best of our knowledge, a new state-of-the-art result on Peptides-func. The results on Peptides-struct are also competitive, with BuNN-Hop remaining within a standard deviation from the best result. The strong performance provides further evidence of the long-range capabilities of BuNNs.
@@ 7 Conclusion
In this work, we proposed Bundle Neural Networks – a new type of GNN that operates via message diffusion on graphs. We gave a formal analysis of BuNNs showing that message diffusion can mitigate issues such as over-smoothing - since the heat equation over vector bundles admits a richer set of fixed points - and over-squashing - since BuNNs can operate at a larger scale than standard MPNNs. We also prove compact uniform approximation of BuNNs, characterizing their expressive power and establishing their superiority over MPNNs. Finally, we showed that BuNNs perform well on a range of synthetic and real-world benchmarks including both inductive and transductive tasks.
@@ Appendix A Limitation and Future work
A limitation of our framework is that while message diffusion allows to operate on different scales of the graph, the computation of the heat kernel for large $t$ requires spectral methods and is therefore computationally expensive. An exciting research direction consists of using existing computational methods to approximate the heat kernel efficiently for large values of $t$. A limitation of our experiments is that we consider inductive graph regression/classification and transductive node regression/classification tasks but no link prediction task. A limitation in our theory is that Theorem  ef thm:unifapp assumes injective positional encodings at the node level, which might only sometimes be available; future work could characterize the expressiveness when these are unavailable.
Appendix B Societal Impact
This paper presents work whose goal is to advance the field of Machine Learning. Our work proposes a novel framework and model that can potentially be used for positive and negative societal impacts. However, due to the technical nature of our work, we are not aware of any particular potential positive or negative societal impact that should be highlighted here.
Appendix C Proofs
@@
Since $\mathbf{\mathcal{L}}_{\mathcal{B}}=\mathbf{O}_{\mathcal{B}}^{T}\mathbf{\mathcal{L}}\mathbf{O}_{\mathcal{B}}$ we get $\mathcal{H}_{\mathcal{B}}(t,u,v)=\mathbf{O}_{\mathcal{B}}^{T}\mathcal{H}(t,u,v)\mathbf{O}_{\mathcal{B}}$. ∎
C.1 Over-smoothing: proofs of Section 4.1.
In this section, we prove the result of over-smoothing. This result follows straightforwardly from Lemma 3.1.
@@
Consider a stable signal $\mathbf{X}\in\ker\mathcal{L}_{\mathcal{B}}$ and pick $u,v\in\mathsf{V}$ such that $\mathbf{O}_{u}\neq\mathbf{O}_{v}$. As $\mathbf{X}$ is stable, it must have 00 Dirichlet energy, so we must have that $\mathbf{O}_{u}\mathbf{y}_{u}=\mathbf{O}_{v}\mathbf{y}_{v}$, but as $\mathbf{O}_{u}\neq\mathbf{O}_{v}$ we have that $\mathbf{y}_{u}\neq\mathbf{y}_{v}$, which holds except in degenerate cases such as when the matrices $\mathbf{O}_{u}$ are reducible, or when the original signal is the zero vector. ∎
C.2 Over-squashing: proofs of Section 4.2.
@@
Let BuNN be a linear layer defined by Equations 23. Then, for any connected graph and nodes $u,\ v$, we have
@@
The result follows from the closed-form solution of the heat kernel from Lemma 3.1. We start by applying the bundle encoder from Equation 2 that updates each node representation as $\mathbf{h}_{v}=\mathbf{O}_{v}^{T}\mathbf{W}\mathbf{O}_{v}\mathbf{x}_{v}+\mathbf{b}$. We therefore compute
@@
Consider n nodes $u,\ v,\ \text{and}\ w_{i},\ \text{for}\ i=1,\ldots n-2$ of a connected graph $\mathsf{G}$ with $2$ dimensional features such that $\mathbf{x}_{v}\neq\mathbf{x}_{w_{i}}$ and $\mathbf{x}_{u}\neq\mathbf{x}_{w_{i}}\ \forall i$. Assume the bundle structure is learned with a node-level MLP. Then, in a given channel, node $v$ can learn to ignore the information from all $w_{i}s$ while keeping information from $u$.
@@
Fix $\mathbf{O}_{v}=\mathbf{O}_{u}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\mathbf{O}_{w_{i}}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. Such maps can always be learned by an MLP, by the assumptions on $\mathbf{x}_{v}$, $\mathbf{x}_{u}$, and $\mathbf{x}_{w_{i}}$ and by the universality of MLPs. Let the weight matrix be $\mathbf{W}=\begin{pmatrix}w_{11}&w_{12}\\ w_{21}&w_{22}\end{pmatrix}$. By Lemma 4.2 we get $\frac{\partial\mathbf{y}_{v}}{\partial\mathbf{x}_{u}}=\mathcal{H}(t,v,u)\mathbf{O}_{v}^{T}\mathbf{W}\mathbf{O}_{u}=\mathcal{H}(t,v,u)\begin{pmatrix}w_{22}&w_{12}\\ w_{21}&w_{11}\end{pmatrix}$ and $\frac{\partial\mathbf{y}_{v}}{\partial\mathbf{x}_{w_{i}}}=\mathcal{H}(t,v,w_{i})\mathbf{O}_{v}^{T}\mathbf{W}\mathbf{O}_{w_{i}}=\mathcal{H}(t,v,w_{i})\begin{pmatrix}w_{21}&w_{22}\\ w_{11}&w_{12}\end{pmatrix}$. Setting $w_{21}$ and $w_{22}$ to 00 gives $\frac{\partial\mathbf{y}^{(1)}_{v}}{\partial\mathbf{x}_{w_{i}}}=\begin{pmatrix}0&0\end{pmatrix}$ and $\frac{\partial\mathbf{y}^{(1)}_{v}}{\partial\mathbf{x}_{u}}=\mathcal{H}(t,v,u)\begin{pmatrix}0&w_{12}\end{pmatrix}\neq\mathbf{0}$, as desired. ∎
C.3 Expressivity of BuNNs: proofs of Section 5.
@@
We start by recalling what a linear feature transformation over a family of graphs $\mathcal{G}$ is:
Definition C.1.
@@
In the proof, we use the following Theorem, which adapts classical results about the universality of MLPs.
Theorem C.2.
@@
Classical theorems such as Theorem 1 in [39] allow us to approximate any continuous function over a compact set by composing a linear map $C$, an activation $\sigma$, and an affine map $A\cdot+b$. By assumption, we can implement the linear map, the activation, and the affine map; hence, by composing them, we can approximate any continuous function over the compact set. ∎
We are now ready to prove the paper’s main result: that, given injective positional encodings, BuNNs are compact universal approximators.
@@
Let $\mathcal{G}$ be a set of connected graphs with injective positional encodings. Then there is a $2$-layer deep BuNN with compact uniform approximation over $\mathcal{G}$.
@@
Reducing to linear approximation. It suffices to show that a BuNN layer can approximate any linear feature transformation $L$ because we can apply classical results such as Theorem C.2 to get universal approximation of $2$-layer deep networks with activation. Following Definition 5.1, we aim to show that we can approximate a linear feature transformation $L$ on any compact subset. For this, we fix $\epsilon>0$, the finite subset $\mathcal{K}\subseteq\mathcal{G}$, and compact feature space $K\subseteq\mathbb{R}^{c}$. In fact, we assume that $K=\mathbb{R}^{c}$ since approximating a linear map on any compact feature space is equivalent to approximating it on the whole space because a linear map defined on a neighborhood of the 00 vector can be extended uniquely to the whole vector space. Our goal is therefore to find a parameterization of a single BuNN layer such that for any graph $\mathsf{G}\in\mathcal{K}$ and for any input feature $\mathbf{X}\in\mathbb{R}^{n_{\mathsf{G}}c}$, we have $\|L_{\mathsf{G}}\left(X\right)-\operatorname{BuNN}_{\mathsf{G}}\left(\mathbf{X}\right)\|_{\infty}<\epsilon$. We will show that $L$ can be parameterized exactly. Since $L$ is linear, it suffices to find a linear BuNN layer that satisfies for any $\mathsf{G}\in\mathcal{K}$ and any $\mathbf{X}\in\mathbb{R}^{n_{\mathsf{G}}c}$, $\frac{\partial\left(L\left(\mathbf{X}\right)\right)_{u}}{\partial\mathbf{x}_{v}}=\frac{\partial\left(\operatorname{BuNN}\mathbf{X}\right)_{u}}{\partial\mathbf{x}_{v}}$. By Lemma 4.2, we have $\frac{\partial\operatorname{BuNN}(\mathbf{X})_{u}}{\partial\mathbf{x}_{v}}=\mathcal{H}(t,u,v)\mathbf{O}_{u}\mathbf{W}\mathbf{O}_{v}^{T}$. Hence, since MLPs are universal and the positional encodings are injective, it suffices to find bundle maps $\mathbf{O}:\mathsf{V}\rightarrow O\left(k\right)$ and $\mathbf{W}$ such that $\frac{1}{n_{\mathsf{G}}d_{u}}\sum_{v\in\mathsf{V}}\mathbf{O}_{u}^{T}\mathbf{W}\mathbf{O}_{v}=\frac{\partial\left(LX\right)_{u}}{\partial X_{v}}$ for every $u,v\in\mathsf{V}$.
@@
Defining the encoder and decoder: In order to find such a BuNN, we first need a linear encoder $\operatorname{lift}:\mathbb{R}^{c}\rightarrow\mathbb{R}^{2ck}$ which will be applied at every node before applying a $2ck$ dimensional BuNN layer. The lifting transformation maps each node vector $\mathbf{X}_{u}$ to the concatenation of $k$ vectors $\mathbf{X}_{u}$ interleaved with $k$ vectors $\vec{\mathbf{0}}\in\mathbb{R}^{c}$. This is equivalent to the linear transformation given by left multiplication by $\left(\mathbf{I}_{c\times c},\mathbf{0},\ldots,\mathbf{I}_{c\times c},\mathbf{0}\right)^{T}\in\mathbb{R}^{2ck\times c}$. After the $2ck$ dimensional BuNN network, we will also need a linear decoder $\operatorname{pool}:\mathbb{R}^{2ck}\rightarrow\mathbb{R}^{c}$ applied to every node individually, which is the sum of the $k$ different $c$-dimensional vectors that are at even indices. This is equivalent to left multiplication by the matrix $\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c},\ldots,\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\in\mathbb{R}^{c\times 2ck}$. These two can be seen as a linear encoder and linear decoder, often used in practical GNN implementations. We prove the result by adding the lifting and pooling layers and using the higher dimensional $\widehat{BuNN}$ layer, so $\operatorname{BuNN}=\operatorname{pool}\circ\widehat{\operatorname{BuNN}}\circ\operatorname{lift}$.

### v2

@@ 6.2 Real-world tasks
We evaluate BuNNs on the Long Range Graph Benchmark (Dwivedi et al., 2022) and the heterophilic tasks from Platonov et al. (2023). We provide the implementation details in Appendix E.
Table 3: Results for the heterophilic tasks. Accuracy is reported for roman-empire and amazon-ratings, and ROC AUC is reported for minesweeper, tolokers, and questions. Best results are denoted by bold. Asterisk denotes that some runs ran out of memory on an NVIDIA A10 GPU (24 GB).
@@
BuNN $\mathbf{91.75\pm 0.39}$ $\mathbf{53.74\pm 0.51}$ $\mathbf{98.99\pm 0.16}$ $\mathbf{84.78\pm 0.80}$ $\mathbf{78.75\pm 1.09}$
Heterophilic datasets. As we have shown in Section 4, BuNNs are provably capable of avoiding over-smoothing. It is, therefore, natural to test how BuNN performs on heterophilic graphs where over-smoothing is recognized as an important limitation (e.g. Yan et al. (2022)). We follow their methodology to evaluate BuNN on the $5$ heterophilic tasks proposed in Platonov et al. (2023). We run the models with $10$ different seeds and report the mean and standard deviation of the test accuracy for roman-empire and amazon-ratings, and mean and standard deviation test ROC AUC for minesweeper, tolokers, and questions. We use the classical baselines from Platonov et al. (2023) and NSD, and provide the hyper-parameters in the Appendix (Section F.2).
Results. We report the results in Table 3. BuNN achieves the best score on all tasks, with an average relative improvement of $4.4\%$. Its score on minesweeper is particularly impressive, which is significantly ahead of the rest and for which BuNN solves the task perfectly. We found that the optimal value of $t$ over our grid search varies across datasets, being $1$ for amazon-ratings and $100$ for roman-empire. BuNN consistently outperforms the sheaf-based model NSD by a large margin, which we believe is due to the fact that NSD learns a map for every node-edge pairs making NSD more prone to overfitting, while BuNNs can act as a stronger regularizer. Such strong performance confirms our theory on over-smoothing from Section 4.1 and showcase the strong modeling capacity of BuNN in heterophilic settings.
Table 4: Results for the Peptides-struct, Peptides-func, and PascalVOC-SP tasks from the Long Range Graph Benchmark (results are ×$\times 100$ for clarity). The best result is bold.
Peptides-func Peptides-struct PascalVOC-SP
Model Test AP $\uparrow$ Test MAE $\downarrow$ Test F1$\uparrow$
GCN $68.60\pm 0.50$ $24.60\pm 0.07$ $20.78\pm 0.31$
GINE $66.21\pm 0.67$ $24.73\pm 0.17$ $27.18\pm 0.54$
GatedGCN $67.65\pm 0.47$ $24.77\pm 0.09$ $38.80\pm 0.40$
DReW $71.50\pm 0.44$ $25.36\pm 0.15$ $33.14\pm 0.24$
SAN $64.39\pm 0.75$ $25.45\pm 0.12$ $32.30\pm 0.39$
GPS $65.34\pm 0.91$ $25.09\pm 0.14$ $\mathbf{44.40\pm 0.65}$
@@
Exphormer $65.27\pm 0.43$ $24.81\pm 0.07$ $39.75\pm 0.37$
BuNN $\mathbf{72.76\pm 0.65}$ $24.63\pm 0.12$ $40.49\pm 0.46$
Long Range Graph Benchmark. In Section 4.2, we showed that BuNNs have desirable properties when it comes to over-squashing and modeling long-range interactions. To verify such claims empirically, we evaluate BuNN on tasks from the Long Range Graph Benchmark (LRGB) (Dwivedi et al., 2022). We consider the Peptides dataset consisting of 15 5351553515\,53515 535 graphs which come with two associated graph-level tasks, Peptides-func and Peptides-struct, and the node classification on the PascalVOC-SP dataset with 11 3551135511\,35511 355 graphs and $5.4$ million nodes where each graph corresponds to an image in Pascal VOC 2011 and each node to a superpixel in that image.
The graph classification task in Peptides-func is to predict the function of the peptide from $10$ classes, while the regression task in Peptides-struct is inferring the 3D properties of the peptides. In both cases, we follow the standard experimental setup detailed by Dwivedi et al. (2022) alongside the updated suggestions from Tönshoff et al. (2023). The performance metric is Average Precision (AP) for Peptides-func and Mean Absolute Error (MAE) for Peptides-struct. The node classification task in PascalVOC-SP consists of predicting a semantic segmentation label for each node out of 21 classes, with performance metric being the macro F1 score. We run each experiment on $4$ distinct seeds and report mean and standard deviation over these runs. Baseline models are taken from Tönshoff et al. (2023) and include MPNNs, transformer models, and the current SOTA models (Gutteridge et al., 2023; He et al., 2023; Rampášek et al., 2022).
Results. We report the results in Table 4. BuNNs achieve, to the best of our knowledge, a new state-of-the-art result on Peptides-func. BuNNs also perform strongly on both other tasks, being second overall on PascalVOC-SP and clearly outperforming all MPNN models, and is third within one standard deviation of the second model on Peptides-struct. The overall strong performance on the LRGB benchmarks confirms our theory on oversquashing from Section 4.2 and provides further evidence of the long-range capabilities of BuNNs.
@@ 7 Conclusion
In this work, we proposed Bundle Neural Networks – a new type of GNN that operates via message diffusion on graphs. We gave a formal analysis of BuNNs showing that message diffusion can mitigate issues such as over-smoothing - since the heat equation over vector bundles admits a richer set of fixed points - and over-squashing - since BuNNs can operate at a larger scale than standard MPNNs. We also prove compact uniform approximation of BuNNs, a first expressivity result of its kind, characterizing their expressive power and establishing their superiority over MPNNs. We then confirmed our theory with carefully designed synthetic experiments. Finally, we showed that BuNNs perform well on heterophilic and long range tasks which are known to be challenging for MPNNs.
8 Acknowledgements
We would like to thank Francesco di Giovanni for insightful conversations at early stages of this work. We are grateful to Teodora Reu for help on early versions of the figures. This research is partially supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1. X.D. acknowledges support from the Oxford-Man Institute of Quantitative Finance and the EPSRC (EP/T023333/1).
@@ Appendix A Limitation and Future work
A limitation of our framework is that while message diffusion allows to operate on different scales of the graph, the computation of the heat kernel for large $t$ requires spectral methods and is therefore computationally expensive. An exciting research direction consists of using existing computational methods to approximate the heat kernel efficiently for large values of $t$. A limitation of our experiments is that we consider inductive graph regression/classification and transductive node regression/classification tasks but no link prediction task. A limitation in our theory is that Theorem 5.3 assumes injective positional encodings at the node level, which might only sometimes be available; future work could characterize the expressiveness when these are unavailable.
Appendix B Proofs
@@
Since $\mathbf{\mathcal{L}}_{\mathcal{B}}=\mathbf{O}_{\mathcal{B}}^{T}\mathbf{\mathcal{L}}\mathbf{O}_{\mathcal{B}}$ we get $\mathcal{H}_{\mathcal{B}}(t,u,v)=\mathbf{O}_{\mathcal{B}}^{T}\mathcal{H}(t,u,v)\mathbf{O}_{\mathcal{B}}$ by the definition of the heat kernel. ∎
B.1 Over-smoothing: proofs of Section 4.1.
In this section, we prove the results on the stable states and on over-smoothing. The first result follows straightforwardly from Lemma 3.1.
@@
Consider a stable signal $\mathbf{Y}\in\ker\mathcal{L}_{\mathcal{B}}$ and pick $u,v\in\mathsf{V}$ such that $\mathbf{O}_{u}\neq\mathbf{O}_{v}$. As $\mathbf{Y}$ is stable, it must have 00 bundle Dirichlet energy, so we must have that $\mathbf{O}_{u}\mathbf{y}_{u}=\mathbf{O}_{v}\mathbf{y}_{v}$, but as $\mathbf{O}_{u}\neq\mathbf{O}_{v}$ we have that $\mathbf{y}_{u}\neq\mathbf{y}_{v}$, which holds except in degenerate cases such as when the matrices $\mathbf{O}_{u}$ are reducible, or when the original signal is the zero vector. ∎
The idea of the second result is that due to the injectivity of the positional encodings on the two nodes, we can set their restriction maps so that the first layer zeroes out the first channel of one node and the second channel of the second. The next $K-1$ layers simply keep this signal fixed by realizing them as a fixed point of the bundle heat equation.
Proposition 4.2.
Let $\mathbf{X}$ be $2$-dimensional features on a connected $\mathsf{G}$, and assume that there are two nodes $u,\ v$ with $\mathbf{p}_{u}\neq\mathbf{p}_{v}$. Then there is an $\delta>0$ such that for every $K$, there exist a $K$-layer deep BuNN with $2$ dimensional stalks, $\phi^{(\ell)}$ is a node-level MLP, ReLU activation, and $\|\mathbf{W}^{(\ell)}\|\leq 1$, with output $\mathbf{Y}$ such that $\|\mathbf{y}_{v}-\mathbf{y}_{u}\|>\delta$.
Proof.
We will show the result for the limit $t\rightarrow\infty$. Let $\mathbf{O}_{w}=\mathbf{Id}$ for $w\neq u$ and $\mathbf{O}_{v}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. These restriction maps can be learned by assumption since $\mathbf{p}_{u}\neq\mathbf{p}_{v}$ and MLPs are universal. Define
$\mathbf{h}=\sum_{w\neq u}\frac{d_{w}}{2|\mathsf{E}|}\mathbf{O}_{w}\mathbf{x}_{w}+\frac{d_{u}}{2|\mathsf{E}|}\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\mathbf{x}_{u}\in\mathbb{R}^{2}$
and denote $h_{0}$ the entry in the first dimension and $h_{1}$ its second entry.
Case 1: $(h_{0})^{2}>0$ or $(h_{1})^{2}>0$.
[…]
Case 2: $(h_{0})^{2}=0$ and $(h_{1})^{2}=0$.
[…]
B.2 Over-squashing: proofs of Section 4.2.
@@
Let BuNN be a linear layer defined by Equations 1, 23 with hyperparameter $t$. Then, for any connected graph and nodes $u,\ v$, we have
@@
The result follows from the closed-form solution of the heat kernel from Lemma 3.1. We start by applying the bundle encoder from Equation 2 that updates each node representation as $\mathbf{h}_{v}=\mathbf{O}_{v}^{T}\mathbf{W}\mathbf{O}_{v}\mathbf{x}_{v}+\mathbf{b}$. Since the $\mathbf{O}_{u}$ do not depend on the signal $\mathbf{X}$, we get
@@
Consider n nodes $u,\ v,\ \text{and}\ w_{i},\text{ for}\ i=1,\ldots n-2$, of a connected graph with $2$ dimensional bundle such that $\mathbf{p}_{v}\neq\mathbf{p}_{w_{i}}$ and $\mathbf{p}_{u}\neq\mathbf{p}_{w_{i}}\ \forall i$. Then in a BuNN layer with MLP $\phi$, at a given channel, the node $v$ can learn to ignore the information from all $w_{i}s$ while keeping information from $u$.
@@
Fix $\mathbf{O}_{v}=\mathbf{O}_{u}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ and $\mathbf{O}_{w_{i}}=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$. Such maps can always be learned by an MLP, by the assumptions on $\mathbf{p}_{v}$, $\mathbf{p}_{u}$, and $\mathbf{p}_{w_{i}}$ and by the universality of MLPs. Let the weight matrix be $\mathbf{W}=\begin{pmatrix}w_{11}&w_{12}\\ w_{21}&w_{22}\end{pmatrix}$. By Lemma 4.3 we get $\frac{\partial\mathbf{y}_{v}}{\partial\mathbf{x}_{u}}=\mathcal{H}(t,v,u)\mathbf{O}_{v}^{T}\mathbf{W}\mathbf{O}_{u}=\mathcal{H}(t,v,u)\begin{pmatrix}w_{22}&w_{12}\\ w_{21}&w_{11}\end{pmatrix}$ and $\frac{\partial\mathbf{y}_{v}}{\partial\mathbf{x}_{w_{i}}}=\mathcal{H}(t,v,w_{i})\mathbf{O}_{v}^{T}\mathbf{W}\mathbf{O}_{w_{i}}=\mathcal{H}(t,v,w_{i})\begin{pmatrix}w_{21}&w_{22}\\ w_{11}&w_{12}\end{pmatrix}$. Setting $w_{21}$ and $w_{22}$ to 00 gives $\frac{\partial\mathbf{y}^{(1)}_{v}}{\partial\mathbf{x}_{w_{i}}}=\begin{pmatrix}0&0\end{pmatrix}$ and $\frac{\partial\mathbf{y}^{(1)}_{v}}{\partial\mathbf{x}_{u}}=\mathcal{H}(t,v,u)\begin{pmatrix}0&w_{12}\end{pmatrix}\neq\mathbf{0}$, as desired. ∎
B.3 Expressivity of BuNNs: proofs of Section 5.
@@
We start by recalling the definition of a linear feature transformation over a family of graphs $\mathcal{G}$:
Definition B.1.
@@
We will need the following Theorem, which adapts classical results on the universality of MLPs.
Theorem B.2.
@@
Classical theorems such as Theorem 1 in (Cybenko, 1989) allow us to approximate any continuous function over a compact set in a finite dimensional vector space by composing a linear map $\mathbf{C}$, an activation $\sigma$, and an affine map $\mathbf{A}\cdot+\mathbf{b}$. Given a finite family of graph $\mathcal{G}$, the space of node features on all graphs is a finite dimensional vector space. By assumption, we can implement the linear map, the activation, and the affine map. Hence, by composing them, we can approximate any continuous function over the compact set. ∎
We are now ready to prove the paper’s main result: that, given injective positional encodings, BuNNs are compact universal approximators of feature transformations.
@@
Let $\mathcal{G}$ be a set of connected graphs with injective positional encodings. Then $2$-layer deep BuNNs with encoder/decoder at each layer and $\phi$ being a $2$-layer MLP have compact uniform approximation over $\mathcal{G}$.
@@
Reducing to linear approximation. It suffices to show that a BuNN layer can approximate any linear feature transformation $L$ because we can apply classical results such as Theorem B.2 to get universal approximation of $2$-layer deep networks with activation. Following Definition 5.1, we aim to show that we can approximate a linear feature transformation $L$ on any compact subset. For this, we fix $\epsilon>0$, the finite subset $\mathcal{K}\subseteq\mathcal{G}$, and compact feature space $K\subseteq\mathbb{R}^{c}$. In fact, we assume that $K=\mathbb{R}^{c}$ since approximating a linear map on any compact feature space is equivalent to approximating it on the whole space because a linear map defined on a neighborhood of the 00 vector can be extended uniquely to the whole vector space. Our goal is therefore to find a parameterization of a single BuNN layer such that for any graph $\mathsf{G}\in\mathcal{K}$ and for any input feature $\mathbf{X}\in\mathbb{R}^{n_{\mathsf{G}}c}$, we have $\|L_{\mathsf{G}}\left(X\right)-\operatorname{BuNN}_{\mathsf{G}}\left(\mathbf{X}\right)\|_{\infty}<\epsilon$. We will show that $L$ can be parameterized exactly. Since $L$ is linear, it suffices to find a linear BuNN layer that satisfies for any $\mathsf{G}\in\mathcal{K}$ and any $\mathbf{X}\in\mathbb{R}^{n_{\mathsf{G}}c}$, $\frac{\partial\left(L\left(\mathbf{X}\right)\right)_{u}}{\partial\mathbf{x}_{v}}=\frac{\partial\left(\operatorname{BuNN}\mathbf{X}\right)_{u}}{\partial\mathbf{x}_{v}}$. By Lemma 4.3, we have $\frac{\partial\operatorname{BuNN}(\mathbf{X})_{u}}{\partial\mathbf{x}_{v}}=\mathcal{H}(t,u,v)\mathbf{O}_{u}\mathbf{W}\mathbf{O}_{v}^{T}$. Hence, since MLPs are universal and the positional encodings are injective, it suffices to find bundle maps $\mathbf{O}:\bigsqcup_{\mathsf{G}\in\mathcal{K}}\mathsf{V}_{\mathsf{G}}\rightarrow O\left(k\right)$ and $\mathbf{W}$ such that $\frac{1}{n_{\mathsf{G}}d_{u}}\sum_{v\in\mathsf{V}}\mathbf{O}_{u}^{T}\mathbf{W}\mathbf{O}_{v}=\frac{\partial\left(LX\right)_{u}}{\partial X_{v}}$ for every $u,v\in\mathsf{G}$ and every $\mathsf{G}\in\mathcal{K}$.
@@
Defining the encoder and decoder: In order to find such a BuNN, we first need a linear encoder $\operatorname{lift}:\mathbb{R}^{c}\rightarrow\mathbb{R}^{2ck}$ which will be applied at every node before applying a $2ck$ dimensional BuNN layer. The lifting transformation maps each node vector $\mathbf{X}_{u}$ to the concatenation of $k$ vectors $\mathbf{X}_{u}$ interleaved with $k$ vectors $\mathbf{0}\in\mathbb{R}^{c}$. This is equivalent to the linear transformation given by left multiplication by $\left(\mathbf{I}_{c\times c},\mathbf{0},\ldots,\mathbf{I}_{c\times c},\mathbf{0}\right)^{T}\in\mathbb{R}^{2ck\times c}$. After the $2ck$ dimensional BuNN network, we will also need a linear decoder $\operatorname{pool}:\mathbb{R}^{2ck}\rightarrow\mathbb{R}^{c}$ applied to every node individually, which is the sum of the $k$ different $c$-dimensional vectors that are at even indices. This is equivalent to left multiplication by the matrix $\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c},\ldots,\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\in\mathbb{R}^{c\times 2ck}$. These two can be seen as a linear encoder and linear decoder, often used in practical GNN implementations. We prove the result by adding the lifting and pooling layers and using the higher dimensional $\widehat{BuNN}$ layer, i.e. we prove that $\operatorname{BuNN}=\operatorname{pool}\circ\widehat{\operatorname{BuNN}}\circ\operatorname{lift}$ can approximate any linear maps which satisfy the encoder and decoder assumption of the Theorem statement.